# Understanding News Thumbnail Representativeness by Counterfactual Text-Guided Contrastive Language-Image Pretraining

### v2

@@ Abstract
This paper delves into the critical challenge of understanding the representativeness of news thumbnail images, which often serve as the first visual engagement for readers when an article is disseminated on social media. We focus on whether a news image represents the main subject discussed in the news text. To serve the challenge, we introduce NewsTT, a manually annotated dataset of news thumbnail image and text pairs. We found that pretrained vision and language models, such as CLIP and BLIP-2, struggle with this task. Since news subjects frequently involve named entities or proper nouns, a pretrained model could not have the ability to match its visual and textual appearances. To fill the gap, we propose CFT-CLIP, a counterfactual text-guided contrastive language-image pretraining framework. We hypothesize that learning to contrast news text with its counterfactual, of which named entities are replaced, can enhance the cross-modal matching ability in the target task. Evaluation experiments using NewsTT show that CFT-CLIP outperforms the pretrained models, such as CLIP and BLIP-2. Our code and data will be made accessible to the public after the paper is accepted.
Understanding News Thumbnail Representativeness by
Counterfactual Text-Guided Contrastive Language-Image Pretraining
Yejun Yoon Soongsil University                        Seunghyun Yoon Adobe Research                        Kunwoo Park Soongsil University
@@ 1 Introduction
Figure 1: An illustration of the proposed method. Given a pair of a news thumbnail image and article, the method generates counterfactual news text and updates a CLIP-like bi-encoder by contrastive learning.
@@
This study aims to investigate the representativeness of news thumbnails, which are images displayed as previews of news articles. Visual content often carries more persuasive impact and leaves a long-lasting impression compared to text Joffe (2008); Newman et al. (2012); Seo (2020). Consequently, the misrepresentation in news thumbnails can cause more critical consequences. Despite its importance, there have been a few studies addressing the representativeness of news images Hessel et al. (2021); Choi et al. (2022). Achieving the goal requires a model to understand the cross-modal relationship between visual and textual content.
@@
Researchers in Natural Language Processing and Computer Vision communities have explored multimodal technologies by addressing general vision and language (V+L) tasks, including visual question answering Antol et al. (2015), image-text retrieval Cao et al. (2022), and visual entailment Xie et al. (2019). Recent approaches have leveraged contrastive learning techniques for representing distinct modalities into a same vector space. CLIP Radford et al. (2021) stands out as one of the pioneering methods that achieve remarkable performance based on contrastive learning, aimed at increasing the similarity between image-text pairs (referred to as positive samples) compared to non-paired image and text (referred to as negative samples). While contrastive learning was not a new concept Oord et al. (2018), CLIP outperformed preexisting methods by utilizing a vast collection of web-based image and text data. Recent advancements have further enhanced the performance by refining learning objectives Li et al. (2022c), incorporating cross-attention layers into the model architecture Li et al. (2021a), and augmenting the training data with high-quality samples Li et al. (2022a); Hao et al. (2023).
@@
Despite the advancements of CLIP-like models in general benchmarks for V+L tasks, we hypothesize that understanding thumbnail representativeness is a still challenging tasks. We highlight two main challenges. (1) Preference for general descriptions in existing datasets: General V+L datasets, such as Conceptual Captions Sharma et al. (2018) or COCO-Caption Chen et al. (2015), were primarily designed to enhance a model’s ability to comprehend general behaviors or scenes. Consequently, these datasets deliberately avoid specific descriptions about entities (e.g., “The man at bat readies to swing at the pitch”). In contrast, news events typically revolve around prominent figures or organizations, leading to news texts that frequently reference named entities Luo et al. (2021). (2) Lack of objective definition and dataset: Representativeness is an inherently abstract and subjective concept, where even humans may disagree with each other. Developing an objective definition is imperative to facilitate the development of proper technologies. A previous study released a dataset of 259 examples Choi et al. (2022), but the samples were labeled according to an abstract definition of representativeness.
@@
Borrowing the principle of five Ws in journalism and news practices The Associate Press (2022), we focus on the representativeness through the lens of Who, on whether an image represents news subjects. Specifically, we ask whether a model can understand the visual manifestation of news subjects, such as showing a politician’s face covered by a news article. Using the manually labelled dataset of 1000 image-text pairs, NewsTT, we evaluated the zero-shot ability of CLIP-like vision and language models by measuring cross-modal similarity scores. To tackle the task more efficiently, this study introduces CFT-CLIP, a CounterFactual Text-guided Contrastive Language-Image Pretraining framework that aligns vision and language representation. As illustrated in Figure 1, the key motivation of the proposed framework is to generate counterfactual news text where the subject of a news event is changed for contrastive updates. In the news text example, the subject is identified as ‘Biden’, and its corresponding token is replaced with ‘Trump’ by the counterfactual text generator. Thus, the generated sentence does not represent the main subject of the news event in the thumbnail image. Our research hypothesis is that learning to contrast the generated counterfactual text with the original news text will enhance the ability of CLIP-like vision language model to assess the representativeness of a thumbnail image.
@@
    We propose a novel task of understanding news thumbnail representativeness with a focus on whether the image represents subjects.
@@
    To serve the task, we introduce NewsTT, a manually annotated dataset of 1,000 news thumbnails and text pairs.
@@
    We propose CFT-CLIP, a counterfactual text guided contrastive language-image pre-training framework.
@@ 2.1 Vision language contrastive pretraining
Researchers have worked on pretraining methods to tackle V+L tasks such as image-text retrieval. Lu et al. (2019) extended BERT to a multi-modal model by co-attentional transformer layers. Another study proposed to learn a universal image and text representation by four pre-training tasks, including masked language modeling and image-text matching Chen et al. (2020). Researchers proposed CLIP, a training scheme of visual representation by learning directly from paired text using contrastive loss Radford et al. (2021). Being trained on a web-scaled dataset, the transformer-based visual representation achieved state-of-the-art results in zero-shot tasks. Another study showed its potential as a reference-free image-captioning metric Hessel et al. (2021). Other researchers presented ALIGN, a bi-encoder trained by a contrastive loss similar to CLIP Jia et al. (2021). While we used CLIP encoders as the target backbone in this study, our method can be applied to a more recent vision-and-language model that incorporates the image-text alignment, such as BLIP-2 Li et al. (2023).
Research has underscored the significance of incorporating hard negative samples into the contrastive objective Nishikawa et al. (2022); Robinson et al. (2021). In response, we introduce a method to generate hard negative text samples using a masked language model. Although a handful of studies have explored the use of masked language models for generating hard negatives Li et al. (2021b); Yao et al. (2022), our approach stands apart due to its distinct token selection and prediction procedures in the step of counterfactual text generation.
@@ 2.2 Multimodal misinformation
This study falls into the broad category of research on multimodal misinformation. Previous research mainly worked on fact verification involving multimodal claims or evidence Zlatkova et al. (2019); Shu et al. (2020); Yao et al. (2023); Rani et al. (2023). In contrast, this study examines the semantic relationship between news thumbnails and text pairs without predicting the veracity of a given image, text, or both. A relevant problem is image repurposing Luo et al. (2021); Abdelnabi et al. (2022), where a real image is used out-of-context to create realistically looking misinformation. To tackle the problem, Luo et al. (2021) proposed NewsCLIPPings, a dataset of out-of-context image detection in the news context. They created the mismatched image-text pairs by swapping the matches of the collected news image and text pairs. Similarly, Mishra et al. (2022) created a dataset of multimodal fact verification, of which false claims were obtained by manipulating the pairs. Unlike the previous research on out-of-context image detection, this study aims to focus on whether news subjects are represented in the thumbnail image. We also created a manually annotated evaluation dataset, contrasting with previous datasets based on synthetic labels Luo et al. (2021); Mishra et al. (2022). Our target task is related to but differs from the manipulated image detection Huh et al. (2018); Rossler et al. (2019), known as deepfakes or cheapfakes, which aimed to detect a fake image generated by AI technologies or simple tools.
3 Target Problem: Understanding News Thumbnail Representativeness
We aim to address the problem of understanding the representativeness of a news thumbnail image for its article. The task is critical due to the two reasons. First, the sharing of news online is abundant Park et al. (2021a); Lee and Ma (2012); Hermida et al. (2012). Second, the thumbnail often serves as the initial and sole visual interaction for social media users when a news link is shared. Visual content is perceived as more credible and leaves a long-lasting impression than text Joffe (2008); Seo (2020); Newman and Zhang (2020). Given these considerations, thumbnails that do not accurately represent their content can misguide the reader’s understanding in online environments.
We formally define the problem as a binary classification. Given a news thumbnail $I$ and its news text $T$, we aim to develop a classifier $f_{\theta}(I,T)$ predicting a binary label $L$ on predicting whether $I$ represents $T$. We assign 1 to $L$ if $I$ represents $T$; otherwise, 0. $I$ is deemed representative of $T$ if the image portrays at least one of the main subjects of a news event. This study aims to develop a vision language model that predicts $L$. We assume the task in a zero-shot setting, where the labeled data is limited. The reference news text $T$ can be either a title or a summary.
@@ 4.1 Raw data collection
NELA-GT-2021 is a headline-oriented [^1] They put the body text whose random tokens are masked. news corpus that encompasses nearly all news articles published by 367 news sources throughout the year 2021 (Gruppi et al., 2022). Since the dataset does not provide web links to thumbnail images, we conducted a supplementary data collection step. By referencing the URLs of news articles available in the dataset, we parsed the HTML document and extracted the thumbnail image URL from the meta tag with the og:image attribute. As a result, we obtained 442,741 pairs of news headlines and thumbnail images, which were sourced from 81 different news media outlets. According to the media ratings provided by Media Bias/Fact Check (MBFC) [^2] https://mediabiasfactcheck.com/, the datasets were well-balanced for the political orientation and trustworthiness ratings. Subsequently, we partitioned this unlabeled dataset into 422,741/10,000/10,000 pairs for train/validation/test purposes while keeping the distribution of the number of articles over different media sources. The validation split was reserved for hyperparameter optimization, and the test split served as the source for labeled data construction.
@@ 4.2 Data annotation
For the annotation process, we sampled 1,000 news articles from the pool of 10,000 instances of the test split. Relying on the MBFC ratings, we sampled 500 articles from the trustworthy outlets, labeled as mixed or high for the trustworthiness rating, and the remaining 500 from the fake news sources, labeled as low. We manually collected the body text for the 1000 articles and ran the OpenAI API for obtaining the summary text by GPT-3.5 Turbo [^3] The used prompt is in Appendix D.. We hired three English-speaking students who frequently read news online from one of the authors’ institutions. Given a news thumbnail, title, and summary text, the annotators were asked to answer whether the news subjects identified in the news text are represented in the thumbnail. To facilitate the annotation process, they were first instructed to specify news subjects identified in the news text, such as named entities or proper nouns, and to select visually displayed ones. We also instructed the annotators to label a thumbnail representative if they think it brings up the image of news subjects without an explicit manifestation. Detailed annotation guidelines can be found in Section E. The annotation results demonstrated a high inter-annotator agreement rate, with a Cronbach’s alpha of 0.863, ensuring the quality and reliability of the labels.
NewsTT, the paired dataset of news thumbnail image and text, consists of 808 representative and 192 non-representative instances. Figure 2 presents an example for each class. The left example shows that the news thumbnail image displays the faces of Biden and Putin, hence labeled as representative. On the other hand, the image on the right does not show the news subject but arouses the potential reader’s interest by showing the surprised faces of the two persons.
Figure 3: An illustration of the counterfactual text generation process. We select named entity tokens in an original text that could indicate news subjects. A BERT-like masked language model generates a counterfactual text by predicting new tokens for the selected entity tokens.
@@ 5 Methods
To tackle the task of understanding thumbnail representativeness in a zero-shot setting where no labeled data is available for training, we employ a thresholding classifier based on embedding similarity,
@@ 5.2 Proposed method: CFT-CLIP
We propose CFT-CLIP, a counterfactual text-guided contrastive language-image pretraining method. The proposed framework aims to improve the CLIP-like bi-encoder by contrastive updates involving the counterfactual text generated from an input text. Figure 3 shows an example of generation. While the original news pertains to an event involving Biden’s summit with the Mexican president, the generated sentence implies a counterfactual event centered on Trump by substituting the token ‘Biden’. This substitution is accomplished via the predictions of a masked language model targeting the named entity token. Based on the news characteristics on the prevalent coverage of named entities Park et al. (2021b); Müller-Budack et al. (2020), we hypothesize that contrasting the original news text with its counterfactual would make the vision and language bi-encoder to obtain a more effective representation for understanding a thumbnail image’s representativeness for the news.

### v3

# Assessing News Thumbnail Representativeness: Counterfactual text can enhance the cross-modal matching ability

@@ Abstract
This paper addresses the critical challenge of assessing the representativeness of news thumbnail images, which often serve as the first visual engagement for readers when an article is disseminated on social media. We focus on whether a news image represents the actors discussed in the news text. To serve the challenge, we introduce NewsTT, a manually annotated dataset of 1000 news thumbnail images and text pairs. We found that the pretrained vision and language models, such as BLIP-2, struggle with this task. Since news subjects frequently involve named entities or proper nouns, the pretrained models could have a limited capability to match news actors’ visual and textual appearances. We hypothesize that learning to contrast news text with its counterfactual, of which named entities are replaced, can enhance the cross-modal matching ability of vision and language models. We propose CFT-CLIP, a contrastive learning framework that updates vision and language bi-encoders according to the hypothesis. We found that our simple method can boost the performance for assessing news thumbnail representativeness, supporting our assumption. Code and data can be accessed at https://github.com/ssu-humane/news-images-acl24.
Assessing News Thumbnail Representativeness:
Counterfactual text can enhance the cross-modal matching ability
[…]
@@ 1 Introduction
Figure 1: An illustration of the key idea of the proposed method. To assess whether a news thumbnail image represents the body text, the method generates counterfactual text to be used as negative samples for contrastive updates.
@@
This study investigates the representativeness of news thumbnails, which are images displayed as previews of news articles. Visual content often carries a more persuasive impact and leaves a long-lasting impression compared to text Joffe (2008); Newman et al. (2012); Seo (2020). Consequently, the misrepresentation in news thumbnails can cause more critical consequences. Despite its importance, there have been a few studies addressing the representativeness of news images Hessel et al. (2021); Choi et al. (2022). Achieving the goal requires a model to understand the cross-modal relationship between visual and textual content.
@@
Researchers in Natural Language Processing and Computer Vision communities have explored multimodal technologies by addressing general vision and language (V+L) tasks, including visual question answering Antol et al. (2015), image-text retrieval Cao et al. (2022), and visual entailment Xie et al. (2019). Recent approaches have leveraged contrastive learning techniques for representing distinct modalities in the same vector space. CLIP Radford et al. (2021) stands out as one of the pioneering methods that achieve remarkable performance, aimed at increasing the similarity between image-text pairs (referred to as positive samples) compared to non-paired image and text (referred to as negative samples). While contrastive learning was not a new concept Oord et al. (2018), CLIP outperformed preexisting methods by utilizing a vast collection of web-based image and text data for pretraining. Recent advancements have further enhanced the performance by refining learning objectives Li et al. (2022b), incorporating cross-attention layers into the model architecture Li et al. (2021a), and augmenting the training data with high-quality samples Li et al. (2022a); Hao et al. (2023).
@@
Despite the progress made in multimodal methods for general vision-and-language benchmarks, we hypothesize that evaluating the representativeness of thumbnails remains difficult, even for the most advanced models, due to two primary challenges. (1) Preference for general descriptions in popular benchmarks: General V+L datasets, such as Conceptual Captions Sharma et al. (2018) or COCO-Caption Chen et al. (2015), were primarily designed to enhance a model’s ability to comprehend general behaviors or scenes. Consequently, these datasets deliberately avoid specific descriptions about entities (e.g., “The man at bat readies to swing at the pitch”). In contrast, news events typically revolve around political actors or organizations, leading to news texts that frequently reference named entities Luo et al. (2021). (2) Lack of dataset labeled according to an objective definition: Representativeness is an inherently abstract and subjective concept, where even humans may disagree with each other. Developing an objective definition is imperative to obtain high-quality labels that different human annotators can agree on and facilitate the development of proper technologies. While there was a human-labeled dataset Choi et al. (2022), they asked for subjective opinions of human annotators without providing objective definitions grounded by journalism principles.
@@
To address (1) and (2), borrowing the principle of five Ws in journalism The Associate Press (2022), we introduce the evaluation dataset of news thumbnails and text pairs, NewsTT. The annotators were instructed to label the 1000 pairs through the lens of Who, on whether an image represents news actors. Using the high-quality labels reflecting the objective definition of representativeness, we evaluated the zero-shot ability of vision and language models to understand news actors’ visual portrayals. We found that CLIP performed better at the task than more recent methods, such as BLIP-2, suggesting the efficiencies of its dual encoder architecture and contrastive objective. To improve its matching ability, this study introduces CFT-CLIP, a counterfactual text-guided contrastive language-image pretraining framework. As illustrated in Figure 1, the key idea of the proposed framework is to generate counterfactual news text where the actor of a news event is changed to be used as hard negatives for contrastive learning. In the news text example, the actor is identified as ‘Biden,’ and its corresponding token is replaced with ‘Trump’ by the counterfactual text generator. Thus, the generated sentence no longer represents the key actor of the news event depicted in the thumbnail image. Our research hypothesis is that learning to contrast the generated counterfactual text with the original news text will enhance the ability of vision language models to assess the representativeness of a thumbnail image.
@@
    We address the problem of assessing news thumbnail representativeness by determining whether a given news image depicts the actors of the news events (Who) according to the principle of 5Ws.
@@
    To serve the task, we introduce NewsTT, a manually annotated dataset of 1,000 news thumbnails and text pairs with high-quality labels. We found that general vision language models struggle with it.
@@
    We propose CFT-CLIP, a counterfactual text-guided contrastive language-image pretraining framework. We found that our simple method could outperform larger pretrained and domain-adapted models, supporting the research hypothesis.
@@ 2.1 Vision language contrastive pretraining
Researchers have worked on pretraining methods to tackle V+L tasks such as image-text retrieval. Lu et al. (2019) extended BERT to a multi-modal model by co-attentional transformer layers. Another study proposed to learn a universal image and text representation by four pre-training tasks, including masked language modeling and image-text matching Chen et al. (2020). Researchers proposed CLIP, a training scheme of visual representation by learning directly from paired text using contrastive loss Radford et al. (2021). Being trained on a web-scaled dataset, the transformer-based visual representation achieved state-of-the-art results in zero-shot tasks. Another study showed its potential as a reference-free image-captioning metric Hessel et al. (2021). Other researchers presented ALIGN, a bi-encoder trained by a contrastive loss similar to CLIP Jia et al. (2021). While we used CLIP encoders as the target backbone in this study, our contrastive learning framework can be applied to a more recent vision-and-language model that incorporates the image-text alignment, such as BLIP-2 Li et al. (2023).
Research has underscored the significance of incorporating hard negative samples into the contrastive objective Nishikawa et al. (2022); Robinson et al. (2021). In response, we introduce a method to generate hard negative text samples using a masked language model. Although a handful of studies have explored the use of masked language models for generating hard negatives Li et al. (2021b); Yao et al. (2022), our approach stands apart due to its distinct token selection and prediction procedures in the step of counterfactual text generation. We compared its effectiveness against the autoregressive generation of GPT-3.5 Turbo in the ablation experiments.
@@ 2.2 Multimodal misinformation
This study falls into the broad category of research on multimodal misinformation. Previous research mainly worked on fact verification involving multimodal claims or evidence Zlatkova et al. (2019); Shu et al. (2020); Yao et al. (2023); Rani et al. (2023). In contrast, this study examines whether news thumbnails and text present the same actor. A relevant problem is image repurposing Luo et al. (2021); Abdelnabi et al. (2022), where a real image is used out-of-context to create realistically looking misinformation. To tackle the problem, Luo et al. (2021) proposed NewsCLIPPings, a dataset of out-of-context image detection in the news context. They created the mismatched image-text pairs by swapping the matches of the collected news image and text pairs. Similarly, Mishra et al. (2022) created a dataset of multimodal fact verification, of which false claims were obtained by manipulating the pairs. Unlike the previous research on out-of-context image detection, this study aims to focus on whether news subjects are represented in the thumbnail image. We also created a manually annotated evaluation dataset, contrasting with previous datasets based on synthetic labels Luo et al. (2021); Mishra et al. (2022). Our target task is related to but differs from the manipulated image detection Huh et al. (2018); Rossler et al. (2019), known as deepfakes or cheapfakes, which aimed to detect a fake image generated by AI technologies or simple tools.
3 Target Problem: Assessing News Thumbnail Representativeness
We aim to address the problem of automatically evaluating the representativeness of a news thumbnail image for its article. The task is critical due to the two reasons. First, the sharing of news online is abundant Park et al. (2021a); Lee and Ma (2012); Hermida et al. (2012). Second, the thumbnail often serves as the initial and sole visual interaction for social media users when a news link is shared. Visual content is perceived as more credible and leaves a long-lasting impression than text Joffe (2008); Seo (2020); Newman and Zhang (2020). Given these considerations, thumbnail images that do not accurately represent their content can misguide the reader’s understanding in online environments.
We formally define the problem as a binary classification. Given a news thumbnail $I$ and its news text $T$, we aim to develop a classifier $f_{\theta}(I,T)$ predicting a binary label $L$ on predicting whether $I$ represents $T$. We assign 1 to $L$ if $I$ represents $T$; otherwise, 0. $I$ is deemed representative of $T$ if $I$ portrays at least one of the actors of a news event, which can be identified from $T$. This study aims to develop a vision language model that predicts $L$ from a pair of $I$ and $T$. We assume the task in a zero-shot setting, where the labeled data is limited. The reference news text $T$ can be either a title or a summary.
@@ 4.1 Raw data collection
NELA-GT-2021 is a headline-oriented [^1] They put the body text whose random tokens are masked. news corpus that encompasses nearly all news articles published by 367 news sources throughout the year 2021 (Gruppi et al., 2022). Since the dataset does not provide web links to thumbnail images, we conducted a supplementary data collection step. By referencing the URLs of news articles available in the dataset, we parsed the HTML document and extracted the thumbnail image URL from the meta tag with the og:image attribute. As a result, we obtained 442,741 pairs of news headlines and thumbnail images, which were sourced from 81 different news media outlets. According to the media ratings provided by Media Bias/Fact Check (MBFC) [^2] https://mediabiasfactcheck.com/, the datasets were well-balanced for the political orientation and trustworthiness ratings. Subsequently, we partitioned this unlabeled dataset into 427,741/5,000/10,000 pairs for train/validation/test purposes while keeping the distribution of the number of articles over different media sources. The validation split was reserved for hyperparameter optimization, and the test split served as the source for labeled data construction.
@@ 4.2 Data annotation
For the annotation process, we sampled 1,000 news articles from the pool of 10,000 instances of the test split. Relying on the MBFC ratings, we sampled 500 articles from the trustworthy outlets, labeled as mixed or high for the trustworthiness rating, and the remaining 500 from the fake news sources, labeled as low. We manually collected the body text for the 1000 articles and ran the OpenAI API for obtaining the summary text by GPT-3.5 Turbo [^3] The used prompt is in Appendix E.. We hired three English-speaking students who frequently read news online from one of the authors’ institutions. Given a news thumbnail, title, and summary text, the annotators were asked to answer whether the news actors identified in the news text are visually portrayed in the thumbnail image. To facilitate the annotation process, we instructed them to identify the actors of the news event by referring to the text and, in turn, to find visually depicted ones. Detailed annotation guidelines can be found in Appendix F. The annotation results demonstrated a high inter-annotator agreement rate, with a Krippendorff’s alpha of 0.705, ensuring the quality and reliability of the labels.
Figure 3: An illustration of the counterfactual text generation process by CFT-CLIP. We select named entity tokens in an original text that could indicate news subjects. A masked language model generates a counterfactual text by predicting new tokens for the selected entity tokens.
Represent (1) Not represent (0)
Title Summary Title Summary
Words 14.9 39.1 14.7 39
Nouns 3.45 8.63 3.75 9.7
Verbs 1.81 4.06 1.68 4.09
Adjectives 0.93 2.39 1.19 2.73
Named entities 2 3.41 1.64 2.76
Table 1: Mean counts of words, part-of-speech units, and named entities measured on the labeled text.
4.3 Data analysis
NewsTT consists of 817 representative and 183 non-representative image-text pairs. Table 1 presents the descriptive characteristics of the dataset for each class. While the samples in the two classes are similarly long, the news text of label 1 tends to include more named entities. Figure 2 presents an example for each class. The label 1 example shows the image portraying Biden and Putin, who are the actors of a news event. The thumbnail image on the right does not show the news actor but arouses the potential reader’s interest by showing the surprised faces.
@@ 5 Methods
To tackle the task of assessing thumbnail representativeness in a zero-shot setting where no labeled data is available for training, we use a thresholding classifier based on the cross-modal similarity,
@@ 5.2 Proposed method: CFT-CLIP
We propose CFT-CLIP, a counterfactual text-guided contrastive language-image pretraining method. The proposed framework aims to improve the vision and language bi-encoder by contrastive updates involving the counterfactual text generated from an input text. Figure 3 shows an example of generation. While the original news pertains to an event involving Biden’s summit with the Mexican president, the generated sentence implies a counterfactual event centered on Trump by substituting the token ‘Biden’. This substitution is accomplished via the prediction of a masked language model targeting the named entity token. Based on the news characteristics on the prevalent coverage of named entities Park et al. (2021b); Müller-Budack et al. (2020), we hypothesize that contrasting the original news text with its counterfactual would make the vision and language bi-encoder to obtain a more effective representation for understanding a thumbnail image’s representativeness.